RFBG Commissioning with Adaptive Secondary - 22 March 2010 UT

Night 3: Working on the Shell Shape, Collimation and Transformations

Observer: JHill, JStorm (LBTO) plus ARakich, DMiller (Tucson)
Telescope Operator: WWack
Software:
Instrument Support::
Telescope Support: MMidkiff
Adaptive Support:: LFini, MXompero (LBTO)

Summary

First part of night had thin cirrus and 1 arcsec seeing. Closed for thick cirrus at 05:30. This allowed time to get caught up on the various wiki logs.

Patched GCS can now run Active Optics as expected.

Doug completed the initial calibration of the pupil position parameters on the WFS.

Reopened at midnight.

Andrew took 3 hours of field aberration data on BS9152.

Did a couple AOSR command tests.

Started a second round of transformation data on StoneL, but it was interrupted by clouds.

Preparations

Torsten sent us a patched version of GCS = V2.2.12B1

Butch thinks he found the closed valve that was causing the winter chiller to have problems.

Jesper processed the initial transformation from last night's StoneL data.

Marco reports that M2 held the flattened shape for the whole time we were sleeping.

John got rid of the old Firefox configuration files for user LBTO. They are stored in /home/LBTO/.mozilla_old_20100322

Details

Open

01:45 Open to sky full of high cirrus. T=+3.3 degC D=-22.7 degC

01:50 pointto Pollux gs=0 TelM="TRACK"
   irc PresetTelescope 0. POSITION TRACK right
Resmsg[0]  = PresetTelescope result status: Error
Resmsg[1]  = OK
Resmsg[2]  = parkActiveProbe: parkProbe() command failed.
Resmsg[3]  = AGW2: Requested position is Out of Range
Is this because of the new transform?

Changed park_y from 420 to 425.0 mm, changed onAxis_y from 612.2 to 612.5 in AIP_IRTC_R.cfg

02:01 Stop and restart GCSR.

02:02 pointto Pollux gs=0 TelM="TRACK" - Adjust IE=-40 CA=-80 at EL=76 AZ=105.

02:08 Hexapod is stuck. Causes OSS to hang. Marco says that it was power cycled today.

02:10 Kill and restart OSS

02:11 Press init button on OSS M2 GUI, but nothing happens. Fortunately, M2 is stuck in a good position.

02:21 Kill and restart OSS a second time. Now it recognizes that hexapod is not homed. Open brakes, Init, Home. Then it is OK. Added a note to Issue 2620.

Pupil looks like last night. Focus is at M2 Zglobal 1.77 mm. Seeing is about 1 arcsec. We have about 1 mag of fairly uniform cirrus.

Marco testing M2

Trying the new GCS patch

02:35 pointto M5_0269 gs=0 TelM="ACQUIRE" PAM- - Adjust CA=-77 at EL=77 deg.

02:40 pointto M5_0269 gs=2 TelM="GUIDE" PA=INDEF=210.37 X=0 Y=173.75

Created /OldRepository/AGW_Data/100322/

02:45 agw_collimate_loop - pupil doesn't look to be in the right position for wfsc 0001, but maybe it was some coma as the next one was OK. Seeing on the guider image is 0.8 arcsec FWHM. Stopped after wfsc 0006

02:52 pointto M5_0269 gs=2 TelM="ACTIVE" - GCSR is processing the WFS images now, but none of us had yet updated the pupil position parameters from last night.

Doug Checking Zernike Signs versus Rotation in IDL

Hand the telescope to Doug while I am updating the parameters for GCS.

Updated these pupil position parameters in AIP_IRTC_R.cfg to match what Doug measured last night. (adjusted again below)
WFScam_hotspot_x        374.0           double  # ccd x coordinate of pixel that represents leftmost x-center of WFS spots
WFScam_hotspot_y        257.0           double  # ccd y coordinate of pixel that represents leftmost y-center of WFS spots
WFS_lenslet_edge_x      354.7           double  # defines the edge of the virtual lenslet grid in x
WFS_lenslet_edge_y      237.0           double  # defines the edge of the virtual lenslet grid in y

03:16 Telescope loses guide lock when Az/El goes off source at 227 73. We don't really know why.

03:22 pointto M5_0269 gs=2 TelM="GUIDE" - failed without star

Now we are back in the silly postage stamp mode of the acquisition image.

03:24 Stop and restart GCSR

03:25 pointto M5_0269 gs=2 TelM="GUIDE" - size of acq image is OK, but no star.

03:28 M5_0269 TRACK Adjust IE=-48 CA=-92 - Did the previous "event" change the pointing?

03:32 pointto M5_0269 gs=2 TelM="GUIDE" - OK

03:33 pointto M5_0269 gs=2 TelM="ACTIVE"

03:34 stopWFS so Doug can continue with IDL.

Made a link from Doug's dd file for takepic.

03:59 pointto M5_0269 gs=2 TelM="ACTIVE" - didn't find star and/or did acq offset on noise

04:02 M5_0269 TRACK Adjust IE=-50 CA=-105

04:04 pointto M5_0269 gs=2 TelM="ACTIVE" -

Running gcs_wfsc_ thru wfsc 000026

04:19 pointto M5_0269 gs=0 TelM="ACTIVE"

Preset hangs collimating M2. OSS was hard to kill, but restarting it restored communication.

04:24 pointto M5_0269 gs=0 TelM="ACTIVE" - but the M5 star is too bright for guiding.

04:26 pointto M5_0269 gs=2 TelM="ACTIVE"

Doug refines the grid position:
04:32 AIP_IRTC_R.cfg:WFS_lenslet_edge_y 223.8 double # defines the edge of the virtual lenslet grid in y

pointto M5_0269 gs=2 TelM="GUIDE" PA=180.37 - didn't find star

pointto M5_0269 gs=2 TelM="ACQUIRE" PA=180.37

Adjust IA=-362 IE=-48 CA=-77 at EL=57 (The trend of IA with warming temperatures continues.)

04:41 pointto M5_0269 gs=2 TelM="GUIDE" PA=180.37 - rotator is not movable

04:42 pointto M5_0269 gs=2 TelM="GUIDE" PA=180.37 -

04:44 pointto M5_0269 gs=2 TelM="GUIDE" PA=170.37 - failed
   irc PresetTelescope 170.37 POSITION GUIDE right
Resmsg[0]  = PresetTelescope result status: Error
Resmsg[1]  = OK
Resmsg[2]  = basicList deserialize error

04:xx pointto M5_0269 gs=2 TelM="GUIDE" PA=170.37 - failed again with different message
irc PresetTelescope 170.37 POSITION GUIDE right
Resmsg[0]  = PresetTelescope result status: Error
Resmsg[1]  = OK
Resmsg[2]  = ERROR: No such handle: -1

04:45 Stop and restart GCSR. See Issue 2629.

04:46 pointto M5_0269 gs=2 TelM="GUIDE" PA=170.37 -

04:47 pointto M5_0269 gs=2 TelM="GUIDE" PA=150.37 X=150.333 Y=86.888

Doug is checking pupil locations with IDL.

05:00 Winter Chiller LWT is -20 degC. Perhaps better, but not fully healthy.

05:02 pointto M5_0269 gs=2 TelM="GUIDE" PA=270.37 X=-150.406 Y=86.724

Doug's verified zernike factors: (different from what he thought last night)
 wfsc_rot_offset = 0.0

zernike_factor = [0.0, 0.0, -1.5, 1.7, -1.6, -1.4, 1.4, 1.7, -1.7, -1.3, $
                  0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, $
                  0.0, 0.0, 0.0, -3.2/2.0]
GCS uses a Z22 value 4x lower than that.

Seeing is still around 1- 1.5 arcsec, but we are up to 2-3 mags of cirrus. Andrew reports clearing over Tucson.

WFScam_hotspot_x 378.0 double # ccd x coordinate of pixel that represents leftmost x-center of WFS spots

That's the last of Doug's initial parameter adjustment.

05:17 Stop and restart GCSR to get parameters.

Active Optics with GCS off-axis

05:19 pointto M5_0269 gs=2 TelM="GUIDE" PA=270.37 X=-150.406 Y=86.724 but clouds are too heavy for this pretty bright guide star. Plenty of scattered moonlight so you can see where the stars fall in the lenslets.

Close

05:30 Close for thickening clouds. We can only barely see the 1st quarter moon through them.

Updating configuration files

The current state of PCSInstrument.conf:
RIGHTPOINTINGMODEL	IRTC_RFBG.20100322.ptmod
RIGHTFOCALLENGTH	123421.4
RIGHTPOINTINGORIGIN	0.0 0.0
RIGHTXYFLIP		false false
RIGHTIAA		0.0
RIGHTPAI		0.0
RIGHTWAVELENGTH		1.400
RIGHTZEROPOINT		-2.400
RIGHTSCALE		1.0

This is the initial pointing model file with a compromise value of IA. /home/telescope/TCS/Configuration/PCS/IRTC_RFBG.20100322.ptmod
LBT Pointing Log File, 2010 Mar 22
T   34   5.4609    0.000   0.0000
  IA       -314.9133    10.40781
  IE        -48.0000     2.56220
  NPAE      -12.6681    12.07098
  CA        -77.0000    14.73604
  AW         -6.6508     1.09723
  AN        +22.8939     1.08815
  TF        +26.4008     3.62162
  HACA2      -4.1062     1.66168
END

Updated the M2 collimation model for last night's on-sky adjustments. /home/telescope/TCS/Configuration/PSF/new/DXSMIRTCFCollimation.dat
[telescope@obs1 new]$ grep Global DXSMIRTCFCollimation.dat
   2.500    -3.630     1.770  -360.000   -120.000     0.000  Global Offsets

Open

Rotator was in velocity mode again.

06:54 Open to thinning clouds T=+2.7 degC D=-18.3 degC.

06:55 pointto M5_1332 gs=0 TelM="ACQUIRE" PAM- Adjust CA=-67 at EL=72

Checking Collimation

06:59 pointto BS9149 gs=0 TelM="ACTIVE" - Seeing is 0.8 arcsec.

07:14 pointto BS9152 gs=0 TelM="ACTIVE"

DX M1 Total X=+0.600 Y=+0.454 Z=+0.203 RX=-8.48 RY=-0.745 EL=61.1 T=2.66

Did I miss a preset here in the log?
Mar 22 07:18:00 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 810.742 0 0 231.948 -72.1328 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:18:40 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 453.069 -422.458 1005.68 38.277 -60.9639 -234.914 234.754 -2.52994 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:19:21 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 70.2776 -172.703 424.591 40.3515 88.1096 -70.0555 23.4147 13.8511 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:20:02 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -14.1801 21.6227 169.087 38.1064 41.4905 8.3865 -38.6508 2.25179 -0 -0 0 -0 0 0 0 0 0 0 -0.256038
Mar 22 07:20:43 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 15.2823 46.635 97.683 -38.9861 -8.14185 61.7829 11.7992 -12.3893 -0 -0 0 -0 0 0 0 0 0 0 2.85481
Mar 22 07:21:23 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 46.0922 49.3827 68.7136 -2.1821 72.3916 44.6147 34.5512 -6.12011 -0 -0 0 -0 -0 0 0 0 0 0 1.01282

above net correction of 1766 nm M1 Z6

07:21 pointto BS9156 gs=0 TelM="ACTIVE" (hot dog shaped image), almost missed the acq image

DX M1 Total X=+0.652 Y=+0.128 Z=+0.691 RX=-7.17 RY=-13.71 EL=45.3 T=+2.66

Mar 22 07:23:53 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -2555.64 0 0 -310.116 75.1501 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:24:34 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -1834.14 0 0 -257.464 -86.5998 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:25:15 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -628.342 241.15 -2636.44 -87.8949 -68.7475 29.8807 -78.211 9.1205 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:25:57 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -341.827 67.0089 -1320.78 29.4008 -100.307 -19.4857 59.4406 41.4055 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:26:39 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 88.1451 -297.221 -409.998 41.0316 -42.5837 -54.4891 25.8846 -68.3083 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:27:20 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 144.82 70.3594 -112.33 45.6397 -7.48216 -45.3219 33.6481 -32.2018 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:28:00 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 108.26 -155.864 8.00074 34.8083 -32.6162 -39.2812 -21.3442 -19.1459 0 0 0 0 0 0 0 0 0 0 0

above net correction of -4472 nm M1 Z6

07:28 pointto BS9149 gs=0 TelM="ACTIVE" (hot dog image again)

DX M1 Total X=+0.799 Y=+0.284 Z=+0.319 RX=-8.51 RY=-6.14 EL=57.1 T=2.67

Mar 22 07:31:49 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 1520.47 0 0 337.323 692.014 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:32:37 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 1136.78 0 0 103.829 -214.905 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:33:17 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 265.174 108.878 2821.11 -63.6979 -104.041 164.886 -67.9927 76.4912 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:33:59 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 157.251 185.332 856.626 -38.1593 -26.5872 103.804 -11.7961 11.3886 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:34:39 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 217.671 71.6371 265.31 -41.0152 -4.31855 71.961 -28.7378 -11.9014 0 0 0 0 0 0 0 0 0 0 0
Mar 22 07:35:19 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 41.8481 -117.132 205.937 36.1558 -36.0109 63.406 5.54415 -16.9043 0 0 0 0 0 0 0 0 0 0 0

above net correction of +4149 nm M1 Z6

07:35 pointto BS9148 gs=0 TelM="ACTIVE" (well out of focus at edge of acq field)

Hit M1 collimation limit before it converged. Active Optics was asking for -2 mm X.
DX M1 Total X=-2.868(red) Y=+0.881(red) Z=-1.099 RX=-20.7 RY=69.6 EL=86.8 T=2.68

07:xx pointto BS9152 gs=0 TelM="ACTIVE"

DX M1 Total X=+0.567 Y=+0.574 Z=+0.037 RX=-9.14 RY=+2.74 EL=64.8 T=2.71

07:51 Stop GCS WFS

Field Aberration Measurements on BS9152
/OldRepository/AGW_Data/100322

07:52 agw_collimate_loop (on-axis) - seeing is 0.6 - 0.7 07:55 GCS says timed out waiting for image.

07:56 Stop and Restart GCSR

07:57 pointto BS9152 gs=0 TelM="GUIDE" (well centered)

GStarSorted ascending wfscimage#
0 047-049
0 050-053
0 missed a couple during operator changeover 61-72
0 77- 79
9 054-056
17 74-76

Do we have a sign error in the calculations of expected aberrations with field now?

18 81-83
0 84- 86
7 88-90
0 91-92
4 94-96
0 97- 101
1 103-105
0 106-108
9 110-112
0 113- 117

8:58 Preset hung at step "GCS sends centroids to PCS" (displayed on IIF)
09:00 Stop and restart GCSR.
09:01 resend preset = OK

17 119-121
0 122-126
18 128-130
0 131-134
7 136- 138

9:25 a bit of cloud getting in the way of things here...
9:28 Adjust *IE=-57 CA=-64*

0 139- 142
4 144-146
0 147-153
1 155-157
0 158- 169
9 170-172

10:06 had to stop and restart GCSR, timing out...problem same as before, rotator has stopped also.

0 173-176
17 177-179
0 180-183
18 184-186
0 187-192
7 194-196
0 197-210
4 212-214
0 215-217
1 219-221
0 222-224

Testing Offload Zernikes command from AOSR

Mon Mar 22 07:03:58.302 2010  SetZernikes started for 22 coefficients to ope DEFAULT from AOSR
Mon Mar 22 07:03:58.353 2010  ApplyZernikes started
Mon Mar 22 07:03:58.403 2010  primary mirror getting Zernike coefficients
Mon Mar 22 07:03:58.455 2010  primary mirror Zernike coefficients received
Mon Mar 22 07:03:58.506 2010  adjusting primary mirror forces
Mon Mar 22 07:03:58.642 2010  primary mirror force adjustments complete
Mon Mar 22 07:03:58.694 2010  secondary mirror getting Zernike coefficients
Mon Mar 22 07:03:58.746 2010  secondary mirror Zernike coefficients received
Mon Mar 22 07:03:58.797 2010  adjusting secondary mirror forces
Mon Mar 22 07:03:58.850 2010  secondary mirror force adjustments complete
Mon Mar 22 07:03:58.902 2010  ApplyZernikes complete
Mon Mar 22 07:03:58.954 2010  SetZernikes complete for 22 coefficient to ope DEFAULT from AOSR

But, this wasn't the new command intended for tip-tilt offload.

10:12 DX M1 TOTAL X=+0.504 Y=+1.012 Z=-0.055 RX=-6.71 RY=+6.54 EL=68.5 T=2.73

Another offload test from AOSR

10:53 pointto StoneLagw ACQUIRE IE=-50 CA=-72

11:00 Restart AOSR, then start another set of test commands.
Mon Mar 22 11:01:42.189 2010  ApplyZernikes started
Mon Mar 22 11:01:42.240 2010  primary mirror getting Zernike coefficients
Mon Mar 22 11:01:42.292 2010  primary mirror Zernike coefficients received
Mon Mar 22 11:01:42.345 2010  adjusting primary mirror forces
Mon Mar 22 11:01:42.485 2010  primary mirror force adjustments complete
Mon Mar 22 11:01:42.538 2010  secondary mirror getting Zernike coefficients
Mon Mar 22 11:01:42.590 2010  secondary mirror Zernike coefficients received
Mon Mar 22 11:01:42.642 2010  adjusting secondary mirror forces
Mon Mar 22 11:01:42.695 2010  secondary mirror force adjustments complete
Mon Mar 22 11:01:42.748 2010  ApplyZernikes complete
Mon Mar 22 11:01:42.801 2010  SetZernikes complete for 22 coefficient to ope DEFAULT from AOSR
Mon Mar 22 11:01:42.854 2010  primary mirror setPointingOffsets started
Mon Mar 22 11:01:42.905 2010  primary mirror adjustMirrorCollimation started with argument 0.0000000e+00
Mon Mar 22 11:01:44.278 2010  primary mirror adjustMirrorCollimation complete with argument 0.0000000e+00
Mon Mar 22 11:01:44.329 2010  primary mirror setPointingOffsets complete
Mon Mar 22 11:01:44.381 2010  secondary mirror setPointingOffsets started
Mon Mar 22 11:01:44.432 2010  secondary mirror adjustMirrorCollimation started with argument 0.0000000e+00
Mon Mar 22 11:01:47.490 2010  secondary mirror adjustMirrorCollimation complete with argument 0.0000000e+00
Mon Mar 22 11:01:47.542 2010  secondary mirror setPointingOffsets complete
01:47.594 2010  AdjustPointing complete for tip 5.0245173e-04, tilt 2.8374785e-04

It seems there are still units issues.

Transform Data on StoneLagw

11:04 pointto StoneLagw gs=0 TelM="ACTIVE" 11:08 takepic 2.0 irtc.20100322.110828.fits (single blank)

11:09 cl < tran_StoneL_315.cl
transtar (31) irtc.20100322.111057.fits cube 47
transtar (21) Preset failed - irtc.20100322.111302.fits cube 48
transtar (19) Preset failed - irtc.20100322.111500.fits cube 49

11:18 Stop and restart GCSR.

11:19 cl < tran_StoneL_315.cl - seeing about 0.8 arcsec, some thin cirrus clouds
transtar (31) irtc.20100322.111944.fits cube 050
transtar (21) irtc.20100322.112220.fits cube 051
transtar (19) irtc.20100322.112456.fits cube 052
transtar (15) irtc.20100322.112735.fits cube 053
transtar (13) irtc.20100322.113017.fits cube 054
transtar (9) irtc.20100322.113256.fits cube 055
transtar (6) irtc.20100322.113535.fits cube 056
transtar (5) irtc.20100322.113813.fits cube 057
transtar (10) irtc.20100322.114103.fits cube 058 (binary star)
transtar (24) irtc.20100322.114338.fits cube 059
transtar (30) irtc.20100322.114616.fits cube 060
transtar (36) wavefront sensing stopped? irtc.20100322.114857.fits cube 061
transtar (12) wavefront sensing not converged irtc.20100322.115153.fits cube 062

I don't understand why WFS stopped working.
Restart IRAF, change date in LBTtools, add keyword file to takeseq
Headers in the above files from takeseq are still wrong as last night.

11:56 tran_StoneL_00.cl (aborted)

11:57 pointto StoneLagw gs=0 TelM="ACTIVE"

12:03 tran_StoneL_00.cl
transtar (35) irtc.20100322.120331.fits cube 063
transtar (31) irtc.20100322.120611.fits cube 064
transtar (21) irtc.20100322.120847.fits cube 065
transtar (19)
transtar (15)
transtar (13) star too faint - unguided irtc.20100322.121651.fits cube 068
transtar (9) heavy clouds irtc.20100322.121925.fits cube 069
aborted by clouds

Close

12:22 Close for heavy clouds T=+2.6 degC

-- JohnHill - 21 Mar 2010
Topic revision: r14 - 30 Mar 2010, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback